OpenStudy (linna):

Find "a" if coficient of line which passes through points P1 (-3,1) and P2(a,a) is 1/a

OpenStudy (linna):

anyone ?

OpenStudy (anonymous):

you're supposed to know this formula: \[\LARGE (y-y_1)=\frac{y_2-y_1}{x_2-x_1} (x-x_1)\] now \[\LARGE k=\frac{y_2-y_1}{x_2-x_1}\] k=1/a substitute... and there you go !